OpenStudy (anonymous):

What is the slope of the line that passes through (2, 3) and (4, 4)?

OpenStudy (anonymous):

U have two points and slope can be calculated by using formula m = y2 - y1/ x2 -x1 Just substitute the values and u get the answer :)

OpenStudy (anonymous):

so the slope equals 2 ?

OpenStudy (anonymous):

4 - 3 / 4 - 2 = 1/2

OpenStudy (anonymous):

Its y2 - y1/x2 - x1

OpenStudy (anonymous):

help me with one more so i can solve it myself ?

OpenStudy (anonymous):

Sure np

OpenStudy (anonymous):

ok so the qeustion is what is the slop of the line taht passes through 1,6 and 3,4 so its going to be set up like 1-3 / 6-4 =0 ?

OpenStudy (anonymous):

It should be the other way x1 = 1 and y1 = 6 x2 = 3 and y2 = 4

OpenStudy (anonymous):

Every Order pair will be of the form (x , y) so first one is x co-ordinate and second one is y co - ordinate

OpenStudy (anonymous):

4 isnt one of the answers

OpenStudy (anonymous):

4 is not the answer dude m = y2 - y1/ x2 - x1 = 4 - 6 / 3 - 1 = -2/2 = -1

OpenStudy (anonymous):

sorry i misinterpreted the directions a little bit, my bad

OpenStudy (anonymous):

Thats ok not a problem :)

OpenStudy (anonymous):

What is the slope of the line that passes through (0, 0) and (4, 1)? so id set it up like 4-0/ 1-0 ? = 3 ?

OpenStudy (anonymous):

No no u again got it wrong always difference of y co - ordinates in numerator and difference of x co-ordinates in the denominator

OpenStudy (anonymous):

1 - 0 / 4 - 0 = 1/4